# A Framework for Quantum Advantage

### v1

@@ IV Elements for Advantage
In what follows, we discuss the necessary elements critical to reach and certify quantum advantage in the problems outlined above. These elements include methods capable of running accurate circuits. Such accuracy can be established through error bars, obtained from error correction, or rigorous error mitigation and detection methods. Finally, we discuss the role that quantum-centric supercomputing infrastructure, and performant quantum hardware will play towards achieving quantum advantage.
@@ IV.1 Methods for running accurate circuits
Error correction Quantum error correction [49] (QEC) refers to a set of techniques designed to protect quantum information from decoherence, gate imperfections, and measurement errors—limitations intrinsic to any practical quantum computing platform. In QEC, information is encoded into logical qubits, fault-tolerant constructs built from multiple physical qubits, such that errors on a subset of physical qubits can be detected and corrected without collapsing the encoded state. QEC lays the groundwork for fault-tolerant quantum computation (FTQC), where quantum algorithms can be performed in a manner that is resilient to noise, provided that the physical error rates remain below a threshold. Among various QEC codes, the surface code [50] has been a leading candidate, for it is particularly suited for 2D architectures like superconducting qubits, and it offers high error thresholds close to 1% per operation [51]. However, this code requires a large overhead—on the order of thousands of physical qubits per logical qubit for realistic noise models—posing a scalability challenge. To address this, quantum low-density parity-check (qLDPC) codes have emerged as a compelling alternative. These codes can offer similar levels of error suppression with a 10x reduction in qubit overhead [52]. This makes them especially attractive for implementing deep-circuit algorithms such as Shor’s algorithm [2] and quantum phase estimation (QPE) [53], which are potential candidates for achieving quantum advantage in areas such as quantum chemistry and integer factorization.
@@
Error mitigation Quantum error mitigation (QEM) refers to the set of techniques that aim to reduce, and sometimes eliminate, the error bias in expectation values caused by noise in quantum circuits [37, 55]. Unlike QEC, QEM techniques use classical post-processing of the noisy output from an ensemble of circuit runs to “mitigate” the effect of noise. Common QEM approaches include readout error mitigation to correct for measurement assignment errors [56, 57], and zero-noise extrapolation (ZNE), in which the noise strength is amplified, with the ultimate goal of “extrapolating” results toward the zero-noise limit [38, 58, 8]. Even in the absence of fault tolerance, some QEM methods can provide accurate estimates of expectation values measured off short-depth quantum circuits, with rigorous error bounds [59, 60]. However, such methods incur significant sampling overhead, often exponential in noise strength, due to the amplification of variance introduced by the quasi-probabilistic weighting. This is the case for probabilistic error cancellation (PEC), an error mitigation method that effectively inverts well-characterized noise channels [60]. The runtime cost for PEC roughly scales as for a -qubit circuit of depth , and hardware two-qubit depolarizing error rates [60]. Despite the exponential scaling, as hardware error rates drop, these methods can provide speed-ups even over classical methods with polynomial runtime scaling [61], at the 100-qubit scale with thousand’s of two-qubit gates. We project that further improvements in hardware error rates will drive the reach of error mitigation to increasingly challenging circuits/observables and circuit depths towards the realization of quantum advantage [62, 63, 64].
The reach of these methods can be further extended with classical HPC, than can further reduce the overhead for error mitigation error mitigation in exchange for classical runtime and potential bias, while leveraging the latest advances in classical simulation. This implies that at a given error rate, one can mitigate upon even larger circuit volumes. We have already seen this using tensor network error mitigation (TEM) [65], shaded light cones for PEC [66], and multi-product formulas (MPF) for algorithmic error mitigation of trotterized time evolution [67]. With these advances, we are confident that we will soon reach a point where quantum methods will outperform existing classical-only methods.
@@
Error detection Lastly, we see promise in post-selected quantum error detection methods. A simple way to view the difference between error correction and error mitigation is to see them as a trade-off between qubits (quantum resources) and sampling (classical resources). Error detection offers a more intermediate goalpost: use extra qubits to detect but not correct errors. This includes error-correcting codes with post-selection, which are also subsumed under error detection. Error detection can have a better (though still exponential) sampling overhead than other error mitigation methods that invert the noise, at the expense of using some extra qubits (though far fewer than full error correction). Post-selected error detection is the process of detecting when a run of the circuit (a ”shot”) has been corrupted by an error, and post-selecting for good shots. The post-selection rate falls exponentially with the circuit volume and the underlying physical noise rate. Unlike other error mitigation approaches, this benefits from an extremely fast sampling speed. In addition to its better sampling overhead, error detection has another appeal, namely that it works on a single-shot basis, not solely on expectation values of observables. Effectively, we slow down the quantum computer but make it produce higher quality samples. This makes error detection compatible with any other error mitigation approach. The combination of better sampling overhead and its single-shot nature makes error detection a promising candidate for achieving finite (but not asymptotic) quantum advantage.
@@ IV.2 Quantum-centric supercomputing infrastructure
A widespread misconception frames quantum advantage as a contest between quantum and classical machines. In practice, however, the relevant benchmark is not quantum systems in isolation, but hybrid quantum–classical systems that combine quantum processing units (QPUs) with classical supercomputers. We expect QPUs to enhance CPUs and GPUs capabilities within a heterogeneous computing model, tackling certain computational problems more efficiently. This emerging paradigm, which we call quantum-centric supercomputing (QCSC in short), positions classical processors as indispensable throughout the quantum computational workflow—handling system control, calibration, input processing such as circuit optimization, and postprocessing of quantum hardware outputs [37, 78]. The objective is therefore not to replace classical computing, but to demonstrate that an integrated hybrid workflow, with quantum computing at its core, can surpass what classical computation alone can achieve.
Realizing near-term quantum advantage, while paving a direct path to large-scale, fault-tolerant machines projected for the latter half of this decade, demands a systematic extension of the HPC architecture so that modular QPUs can be orchestrated as additional resources. At the hardware level, we envision QPUs as integral components of quantum-centric supercomputers—hybrid systems where QPUs act as specialized co-processors integrated with classical compute elements like CPUs and GPUs at different levels of tightness to enable seamless deployment of hybrid quantum-classical algorithms. Pasqal and IBM quantum computers are already being deployed in high-performance computing centers alongside classical resources[9, 79]. And both companies recently announced the release of a SLURM plugin developed with the HPC community aimed at making the integration of QPUs in HPC centers standard and seamless [80].
This set of tools and capabilities will be instrumental in heterogeneous algorithm development through a modular approach in which the different algorithmic pieces can be isolated and benchmarked for performance. The output of this process should be aimed towards reconciling the innate parallelism of modern HPC programming models with the way quantum workloads are defined, optimized, and deployed. Algorithm optimization should leverage workflow management capabilities (e.g. leveraging real-time sampling in variational approaches like SQD for more efficient or faster time-to-solution outcomes).
Extending the tools and infrastructure already familiar to the HPC community is essential to accelerating the adoption of quantum computing across the broader scientific and industrial ecosystem.
IV.3 Performant quantum hardware
@@
Achieving quantum advantage in the short term is fundamentally constrained by the performance of the underlying hardware. In particular, the number of qubits, the achievable algorithmic depth, the fidelity of qubit operations, and the efficiency of quantum–classical orchestration, collectively set the ceiling for any credible claim of advantage. Among the various quantum architectures under development, superconducting quantum processors [81] and neutral atoms [82] stand out as particularly well positioned to reach this milestone, as both have recently surpassed the 100-qubit threshold and demonstrate promising progress across all key performance dimensions. Other platforms, such as trapped-ion quantum processors [83], offer a compelling alternative path to quantum computation. These systems have been deployed in cloud-accessible environments and are already being integrated into early-stage hybrid quantum-classical workflows. While our focus in this work will remain on superconducting circuits and neutral atom systems, the broader landscape of quantum hardware continues to evolve rapidly, with multiple modalities contributing to progress toward quantum advantage.
@@
Superconducting quantum processors are among the most technologically advanced platforms to date. They offer high-fidelity, high-speed gate operations, modularity, and compatibility with integrated control electronics.
IBM’s quantum processors, for example, employ transmon qubits arranged in modular, scalable architectures [84, 85]. The current 156-qubit Heron QPU achieves 250+ s relaxation times, error per layer gate (EPLG) of [85], and supports circuit layer operations per second (CLOPS) above 250,000. [9, 86]. Real-time classical controllers enable dynamic circuits and mid-circuit measurement and feed-forward, allowing for error detection and mitigation protocols.
Execution of hybrid quantum-classical workloads is supported by Qiskit, which provides a cloud-based runtime environment for low-latency quantum-circuit execution with classical feedback [87]. This framework integrates seamlessly in a serverless environment, orchestrates distributed hybrid quantum–classical tasks across containerized back-ends, and provides tools for domain-specific compilation, optimization, and error mitigation routines. Together, these components reduce overhead in iterative algorithms and enables a high-performance development environment that extends the practical reach of near-term devices toward practical demonstrations of quantum advantage.
@@
Neutral-atom quantum computers are rapidly emerging as strong candidates for demonstrating quantum advantage, owing to their scalability, flexible connectivity, and ability to operate in both analog and digital modes [88]. These platforms trap individual atoms in programmable optical tweezers and control their interactions via Rydberg excitation, enabling the realization of both Ising-type analog dynamics and gate-based digital quantum circuits.
@@ V Conclusion and future outlook
We maintain that credible evidence of quantum advantage is likely to emerge in one of the areas highlighted here within the next two years, contingent on sustained coordination between the high-performance-computing and quantum-computing communities, and most likely in the areas we have highlighted here.
Once these claims begin to emerge, they must be assessed on a common basis according to the two criteria articulated in this article. Furthermore, quantum advantage is unlikely to materialize as a single, definitive milestone. Instead, it should be understood as a sequence of increasingly robust demonstrations. Any claim of quantum advantage should be framed as a falsifiable scientific hypothesis, one that invites scrutiny, reproduction, and challenge from the broader community. A cautious iterative and interpretive stance is therefore essential until state-of-the-art classical methods have had adequate opportunity to challenge the claim. However, should a classical approach later supersede a previously reported quantum advantage benchmark, the outcome does not signify failure; rather, it exemplifies constructive behavior that drives progress in both classical and quantum computation.
@@ Acknowledgements.
The authors thank the IBM Quantum and Pasqal teams for fruitful discussions and comments on the manuscript.References
@@
- Eisert et al. [2020] J. Eisert, D. Hangleiter, N. Walk, I. Roth, D. Markham, et al., Nature Reviews Physics 2, 382 (2020).
@@
- Villalonga et al. [2024] B. Villalonga, X. Mi, S. Mandra, A. Bengtsson, P. Klimov, Z. Chen, S. Hong, C. Erickson, I. Drozdov, et al., Nature 634, 328 (2024).
@@
- Georgescu et al. [2014] I. M. Georgescu, S. Ashhab, and F. Nori, Rev. Mod. Phys. 86, 153 (2014).
- Johnson et al. [2014] T. H. Johnson, S. R. Clark, and D. Jaksch, EPJ Quantum Technology 1, 10 (2014).
- Di Meglio et al. [2024] A. Di Meglio, K. Jansen, I. Tavernelli, E. Amico, M. Berta, et al., PRX Quantum 5, 037001 (2024).
- Alexeev et al. [2024] Y. Alexeev, M. Amsler, M. A. Barroca, S. Bassini, T. Battelle, et al., Future Generation Computer Systems 160, 666 (2024).
@@
- Aharonov et al. [2025] D. Aharonov, O. Alberton, I. Arad, Y. Atia, E. Bairey, Z. Brakerski, I. Cohen, O. Golan, I. Gurwich, O. Kenneth, et al., arXiv preprint arXiv:2503.17243 (2025).
- Zimborás et al. [2025] Z. Zimborás, B. Koczor, Z. Holmes, E.-M. Borrelli, A. Gilyén, H.-Y. Huang, Z. Cai, A. Acín, L. Aolita, L. Banchi, et al., arXiv preprint arXiv:2501.05694 (2025).
@@
- McArdle et al. [2020] S. McArdle, S. Endo, A. Aspuru-Guzik, S. C. Benjamin, and X. Yuan, Reviews of Modern Physics 92 (2020).

### v2

@@ IV Elements for Advantage
In what follows, we discuss crucial elements for the realization of a certifiable quantum advantage. These elements involve capabilities that enable the accurate execution of quantum circuits, with accuracy established through rigorous error bars. Finally, we discuss the role that quantum-centric supercomputing infrastructure, and performant quantum processors will play towards the realization of a quantum advantage.
@@ IV.1 Methods for running accurate circuits
Error correction Quantum error correction [50] (QEC) refers to a set of techniques designed to protect quantum information from decoherence, gate imperfections, and measurement errors—limitations intrinsic to any practical quantum computing platform. In QEC, information is encoded into logical qubits: fault-tolerant constructs built from multiple physical qubits, such that errors on a subset of physical qubits can be detected and corrected without collapsing the encoded state. QEC lays the groundwork for fault-tolerant quantum computation (FTQC), where quantum algorithms can be performed in a manner that is resilient to noise, provided that the physical error rates remain below a threshold. Among various QEC codes, the surface code [51] has been a leading candidate, for it is particularly suited for 2D architectures like superconducting qubits, and it offers high error thresholds close to 1% per operation [52]. However, this code requires a large overhead—on the order of thousands of physical qubits per logical qubit for realistic noise models—posing a scalability challenge. To address this, quantum low-density parity-check (qLDPC) codes have emerged as a compelling alternative. These codes can offer similar levels of error suppression with a 10x reduction in qubit overhead [53]. This makes them especially attractive for implementing deep-circuit algorithms such as Shor’s algorithm [2] and quantum phase estimation (QPE) [54], which are potential candidates for achieving quantum advantage in areas such as integer factorization and quantum chemistry.
@@
Error mitigation Quantum error mitigation (QEM) encompasses a suite of techniques aimed at reducing—or in some cases, eliminating—the bias in expectation value estimates caused by noise in quantum circuits [42, 56]. Unlike quantum error correction (QEC), QEM operates through classical post-processing of noisy outputs from ensembles of circuit executions, without requiring additional qubits or fault-tolerant architectures. Common QEM strategies include readout error mitigation, which corrects for measurement assignment errors [57, 58], and zero-noise extrapolation (ZNE), where noise strength is deliberately amplified to enable extrapolation to the zero-noise limit [43, 59, 8]. Even in the absence of fault tolerance, several QEM methods have demonstrated the ability to yield accurate expectation values from short-depth circuits, with rigorous error bounds [60, 61]. However, these methods often incur significant sampling overhead, particularly due to the variance amplification introduced by quasi-probabilistic weighting. This is especially true for probabilistic error cancellation (PEC), which effectively inverts well-characterized noise channels [61]. The runtime cost of PEC scales approximately as for an -qubit circuit of depth and two-qubit depolarizing error rate . Despite this exponential scaling, as hardware error rates continue to decline, QEM techniques can outperform classical methods even when the latter scale polynomially [62] at the 100-qubit scale with thousands of two-qubit gates. We anticipate that continued improvements in hardware fidelity will further extend the reach of error mitigation to deeper circuits and more complex observables, bringing us closer to the realization of quantum advantage from expectation values [63, 64, 65].
The reach of quantum error mitigation methods can be significantly extended through the integration of classical high-performance computing (HPC). By trading off additional classical runtime and tolerating controlled bias, classical resources can reduce the overhead associated with quantum error mitigation, enabling mitigation over larger circuit volumes at fixed hardware error rates. This synergy can leverage the latest advances in classical simulation. Notable examples include tensor network error mitigation (TEM) [13, 66], which employs tensor networks to invert the noise channel; shaded light cone techniques for probabilistic error cancellation (PEC) [67], which limit the scope of noise inversion to causally relevant regions; and multi-product formulas (MPF) for algorithmic error mitigation in trotterized time evolution [68], which improve accuracy without increasing circuit depth. These developments suggest that, even in the absence of full fault tolerance, quantum methods augmented by classical post-processing and simulation can soon surpass the capabilities of classical-only approaches. As hardware continues to improve and hybrid strategies mature, we anticipate that this convergence will play a pivotal role in the realization of quantum advantage.
@@
Error detection Post-selected quantum error detection methods represent a promising intermediate approach between full quantum error correction and classical error mitigation. Conceptually, one can view the distinction between error correction and error mitigation as a trade-off between quantum and classical resources—qubits versus sampling. Error detection occupies a middle ground: it uses additional qubits to detect, but not correct, errors. This includes schemes based on error-correcting codes with post-selection, which fall under the broader umbrella of error detection.
Compared to error mitigation methods like probabilistic error cancellation (PEC), error detection can offer improved (though still exponential) sampling overhead, while requiring significantly fewer qubits than full error correction. Post-selected error detection is the process of detecting when a run of the circuit (a ”shot”) has been corrupted by an error, and post-selecting for good shots. The post-selection rate decreases exponentially with circuit volume and physical noise rate, but this approach benefits from extremely fast sampling speeds.
Another key advantage of error detection is its compatibility with single-shot measurements, rather than relying solely on expectation values. This allows quantum devices to produce higher-quality samples at the cost of reduced throughput—effectively slowing down the quantum computer to improve output fidelity. Moreover, error detection is inherently compatible with other error mitigation techniques, making it a flexible and powerful tool. The combination of better sampling overhead and its single-shot applicability makes error detection a promising candidate for achieving finite (but not asymptotic) quantum advantage.
@@ IV.2 Quantum-centric supercomputing infrastructure
A common misconception frames quantum advantage as a competition between quantum and classical machines. In practice, however, the relevant benchmark is not quantum systems in isolation, but hybrid quantum–classical systems that combine quantum processing units (QPUs) with classical supercomputers. As discussed in earlier sections, HPC can already play a critical role in extending the reach of quantum algorithms run on current pre-fault tolerant hardware —enhancing error mitigation and scaling ground-state simulations.
This emerging paradigm, which we refer to as quantum-centric supercomputing (QCSC), envisions QPUs as specialized co-processors tightly integrated with CPUs and GPUs. Classical processors remain essential throughout the quantum workflow, supporting tasks such as system control, calibration, circuit optimization, and post-processing [9, 66, 78]. The goal is not to replace classical computing, but to demonstrate that a hybrid architecture can outperform classical systems alone.
Realizing near-term quantum advantage—and preparing for fault-tolerant systems anticipated later in this decade—requires extending HPC infrastructure to support modular QPU integration. This vision is already taking shape: Pasqal and IBM quantum systems are being deployed in HPC centers [9, 79], and both companies have released SLURM plugins to streamline QPU integration [80].
These developments support a modular approach to algorithm design, where components can be independently benchmarked and optimized. Aligning quantum workloads with the parallelism of HPC programming models will be key. Workflow management tools—such as real-time sampling in variational algorithms like SQD—can further accelerate convergence and improve time-to-solution.
IV.3 Performant quantum processors
@@
Achieving quantum advantage in the near term is fundamentally constrained by the performance of the underlying hardware. In particular, the number of qubits, the achievable algorithmic depth, the fidelity of qubit operations, and the efficiency of quantum–classical orchestration, collectively set the ceiling for any credible claim of advantage. Among the various quantum architectures under development, superconducting quantum processors [81] and neutral atoms [82] stand out as particularly well positioned to reach this milestone, as both have recently surpassed the 100-qubit threshold and demonstrate promising progress across all key performance dimensions. Other platforms, such as trapped-ion quantum processors [83], also offer a compelling alternative path to quantum computation. These systems have been deployed in cloud-accessible environments and are already being integrated into early-stage hybrid quantum-classical workflows. While our focus in this work will remain on superconducting circuits and neutral atom systems, the broader landscape of quantum hardware continues to evolve rapidly, with multiple modalities contributing to progress toward quantum advantage.
@@
Superconducting quantum processors are among the most technologically advanced platforms to date [84, 85]. They offer high-fidelity, high-speed gate operations, modularity, and compatibility with integrated control electronics.
IBM’s quantum processors, for example, employ transmon qubits arranged in modular, scalable architectures [86, 84]. The current 156-qubit Heron QPU achieves an error per layer gate (EPLG) of [84], and supports circuit layer operations per second (CLOPS) above 250,000 [9, 87]. Real-time classical controllers enable dynamic circuits and mid-circuit measurement and feedforward, allowing for error detection and mitigation protocols. The fast sampling rates of superconducting qubits makes them particularly attractive for near-term explorations of quantum advantage that rely on a large number of samples – such as computations that rely on error mitigation or post-selected error detection capabilities, and algorithms like SQD.
Finally, the hardware is ably supported by a comprehensive software stack, Qiskit, which provides a cloud-based runtime environment for low-latency quantum-circuit execution with classical feedback [88]. This framework integrates seamlessly in a serverless environment, orchestrates distributed hybrid quantum–classical tasks across containerized back-ends, and provides tools for domain-specific compilation, optimization, and error mitigation routines. Together, these components reduce overhead in iterative algorithms and enables a high-performance development environment that extends the practical reach of near-term devices toward practical demonstrations of quantum advantage.
@@
Neutral-atom quantum computers are rapidly emerging as strong candidates for demonstrating quantum advantage, due to their scalability, flexible connectivity, and ability to operate in both analog and digital modes [89, 90]. These platforms trap individual atoms in programmable optical tweezers and control their interactions via Rydberg excitation, enabling the realization of both Ising-type analog dynamics and gate-based digital quantum circuits.
@@ V Conclusion and future outlook
We expect that credible evidence of quantum advantage is likely to emerge in one of the areas highlighted here within the next two years, enabled by sustained coordination between the high-performance-computing and quantum-computing communities.
@@ Acknowledgements.
The authors thank the IBM Quantum and Pasqal teams for fruitful discussions and insightful comments on the manuscript.References
@@
- Abanin et al. [2025] D. A. Abanin, R. Acharya, L. Aghababaie-Beni, G. Aigeldinger, A. Ajoy, et al., arXiv preprint arXiv:2506.10191 (2025).
- Filippov et al. [2023] S. Filippov, M. Leahy, M. A. C. Rossi, and G. García-Pérez, (2023), arXiv:2307.11740 [quant-ph] .
- Haghshenas et al. [2025] R. Haghshenas, E. Chertkov, M. Mills, W. Kadow, S.-H. Lin, et al., (2025), arXiv:2503.20870 .
- Bravyi et al. [2018] S. Bravyi, D. Gosset, and R. König, Science 362, 308 (2018).
@@
- Villalonga et al. [2024] B. Villalonga, X. Mi, S. Mandra, A. Bengtsson, P. Klimov, et al., Nature 634, 328 (2024).
@@
- Aharonov et al. [2025] D. Aharonov, O. Alberton, I. Arad, Y. Atia, E. Bairey, et al., arXiv preprint arXiv:2503.17243 (2025).
- Zimborás et al. [2025] Z. Zimborás, B. Koczor, Z. Holmes, E.-M. Borrelli, A. Gilyén, et al., arXiv:2501.05694 (2025).
@@
- Fuller et al. [2025] B. Fuller, M. C. Tran, D. Lykov, C. Johnson, M. Rossmannek, et al., arXiv preprint arXiv:2502.01897 (2025).
@@
- Acharya et al. [2024] R. Acharya, D. A. Abanin, L. Aghababaie-Beni, I. Aleiner, T. I. Andersen, et al., Nature (2024).
@@
- Wurtz et al. [2023] J. Wurtz, A. Bylinskii, B. Braverman, J. Amato-Grill, S. H. Cantu, et al., (2023), arXiv:2306.11727 [quant-ph] .